# Dissipative solutions to Stochastic 3D Euler equations

### v2

@@ A.1. Estimates for transport equations
We report some well-known estimates regarding smooth solutions $f:\mathbb{R}^{3}\to\mathbb{R}$ of the transport equation
(A.1) $\begin{cases}&\partial_{t}f+v\cdot\nabla f=g,\\ &f(\cdot,t_{0})=f_{0},\end{cases}$
where $v:\mathbb{R}^{3}\to\mathbb{R}^{3}$ is a smooth vector field and $g:\mathbb{R}^{3}\to\mathbb{R}$ is a smooth scalar field. The following proposition is a slight modification of [BuDLSzVi19, Proposition B.1].
@@
(A.5) $\|\nabla\phi(t)-\mathrm{Id}\|_{0}$ $\lesssim|t-t_{0}|\|v_{[t_{0},t]}\|_{1},$
(A.6) $[\phi(t)]_{N}$ $\lesssim|t-t_{0}|[v_{[t_{0},t]}]_{N},\quad\forall N\geq 2.$
@@ A.2. Antidivergence estimates
(A.7) $\mathcal{F}[T_{\mathfrak{m}}(f)](k)=\mathcal{F}[f](k)\mathfrak{m}(k),\quad k\in\mathbb{Z}^{3}.$
@@
Let $v\in C^{\infty}(\mathbb{T}^{3},\mathbb{R}^{3})$ and $l,\epsilon,\lambda$ be positive constants. Let $v_{l}:=\mathbf{P}_{\leq l^{-1}}v$ and consider, for any $m\in\mathbb{Z}$, the solution $\xi_{m}$ to the vector-valued transport equation
@@
Then, for every $N\in\{0,1,2\}$ and every $t\in\mathbb{R}$ we have
@@
with $\epsilon_{n_{0},j}$ being obtained by Lemma˜A.2 applied to the Fourier multiplier operator $\mathbf{P}_{2^{j}}$. Notice that, by construction, $\xi_{m}$ is smooth and space-periodic, since $v_{l}$ is so.
In order to show (A.20), let us consider $(m,k)\in\mathbb{Z}\times(\mathbb{Z}^{3})^{*}$ and $j_{k}\in\mathbb{N}$ such that $2^{j_{k}}\geq\frac{5}{4}\lambda|k|$. Then
(A.23) $\mathbf{P}_{\gtrsim 3\lambda/8}=\sum_{3\lambda/8\leq 2^{i}\leq 2^{j_{k}}}\mathbf{P}_{2^{i}}+\mathbf{P}_{>2^{j_{k}}}=:T_{\mathfrak{m}_{1}}+\sum_{j>j_{k}}T_{\mathfrak{m}_{j}},$
@@
(A.24) $\operatorname{Supp}\,\mathfrak{m}_{1}\subset B(0,2^{j_{k}+1}),\,\quad\mathfrak{m}_{1}\equiv 1\,\text{ on }\,\overline{B(0,2^{j_{k}})}\setminus B(0,\frac{3}{4}\lambda|k|),\quad\mathfrak{m}_{j}\equiv 0\text{ on }\overline{B}(0,\frac{5}{4}\lambda|k|).$
Moreover for every $t\in\operatorname{Supp}_{t}a_{m,k}$, Assumption 1 of this proposition and Eq.˜A.5 imply that
@@
Therefore, Lemma˜A.2 applied to $T_{m_{1}}$, and noticing that $\mathfrak{m}_{1}(\lambda\nabla(k\cdot\xi_{m}))\equiv 1$, $\nabla^{h}\mathfrak{m}_{1}(\lambda\nabla(k\cdot\xi_{m}))\equiv 0$ for every $h\geq 1$, giving:
(A.26) $T_{\mathfrak{m}_{1}}(a_{m,k}e^{i\lambda k\cdot\xi_{m}})$ $-\sum_{3\lambda/8\leq 2^{i}\leq 2^{j_{k}}}\mathcal{\epsilon}_{n_{0},i}(k\cdot\xi_{m},a_{m,k})e^{i\lambda k\cdot\xi_{m}}$
@@
At last, Eq.˜A.20 follows from applying Lemma˜A.2 to $T_{\mathfrak{m}_{j}}$ for all $j>j_{k}$. Indeed,
@@
Now, for fixed $(m,k)$, we want to estimate $\|\epsilon_{n_{0}}^{\lambda}(k\cdot\xi_{m},a_{m,k})\|_{0}$. Based on Eq.˜A.22, we have to consider $\epsilon_{n_{0},j}(k\cdot\xi_{m},a_{m.k})$, which is given by Eq.˜A.9.
Let $Z$ be defined by Eq.˜A.12 and $Z_{0}(r):=r^{-1}Z(r)$, then
@@
Therefore, Assumption 1 and Proposition˜A.1 imply that
@@
Then, by means of Eqs.˜A.13 and A.14, we have that
@@
Then, Eqs.˜A.32 and A.34 imply that
@@
So, Eq.˜A.20 allows to conclude as follows:
@@
(A.41) $\lesssim a_{F}[s_{0}\,\lambda^{N-1}+s_{n_{0}+1}\lambda^{-n_{0}-1}],$
@@
Under the same assumptions of Proposition˜A.3, let $l>\frac{100}{\lambda}$ and $w\in[C^{\infty}(\mathbb{T}^{3})]^{3}$ such that there exists a non-decreasing sequence $(b_{n})_{n\in\mathbb{N}}$ satisfying for every $n,m\in\mathbb{N}$
@@
where $D_{t,l}:=\partial_{t}+v_{l}\cdot\nabla$ and $j\leq(n_{0}+1)\vee 2$, and that for any $n\in\mathbb{N}$
@@
Then for every $N\leq 1$ and $h_{0}\in\mathbb{N}\setminus\{0,1\}$ it holds
(A.46) $\|D_{t}\mathcal{R}F(t)\|_{N}$ $\lesssim\Big[s_{0}(c\lambda)^{-1}+c^{-1}s_{n_{0}+1}\lambda^{-n_{0}-1}+s_{0}\lambda^{-1}\|v\|_{1}$
(A.47) $\quad+(1+\lambda^{-n_{0}}s_{n_{0}+1})(b_{0}+l\|v\|_{1}+\|v\|_{0})+s_{0}(l\lambda)^{-h_{0}}\lambda^{2}\|v\|_{1}\Big]\lambda^{N}a_{F},$
@@
Since $D_{t,l}F=\sum_{k,m}D_{t,l}a_{m,k}e^{i\lambda k\cdot\xi_{m}}$, then the previous Proposition˜A.3 implies that
@@
$\lesssim\lambda^{N+1}\|v\|_{0}\|\sum_{m,k}\epsilon^{\lambda}_{n_{0}}(k\cdot\xi_{m},a_{m,k})\|_{0}$
@@
where the last inequality is due to Eq.˜A.38. Instead, by considering $\mathcal{R}$ as a Fourier multiplier with coefficients $\mathcal{F}(\mathcal{R})(k)$, see [RT25] for an explicit formula, we have that
(A.55) $-I_{1}$ $=\sum_{2^{j}\gtrsim\lambda}\sum_{k,n}i\eta e^{ik\cdot x}\mathcal{F}(v_{l})(k-\eta)\mathcal{F}[F_{j}](\eta)[\mathcal{F}(\mathcal{R})(k)-\mathcal{F}(\mathcal{R})(\eta)]$
$=\sum_{2^{j}\gtrsim\lambda}\sum_{k,n}i\eta\mathcal{F}(v_{l})(k-\eta)\mathcal{F}(F_{j})(\eta)e^{ik\cdot x}\Big[\sum_{h=1}^{h_{0}}\frac{1}{h!}[(k-\eta)\cdot\nabla_{\eta}]^{h}\mathcal{F}(\mathcal{R})(\eta)$
$+\frac{1}{h_{0}!}\int_{0}^{1}[(k-\eta)\cdot\nabla_{\eta}]^{h_{0}+1}\mathcal{F}(\mathcal{R})(\eta+\sigma(k-\eta))(1-\sigma)^{h_{0}}d\sigma\Big]$
$=:J_{1}+J_{2},$
where $F_{j}=\mathbf{P}_{2^{j}}F.$ Now,
(A.56) $\|J_{1}\|_{N}$ $\lesssim\sum_{h=1}^{h_{0}}\|\nabla^{h}v_{l}\nabla\mathcal{R}_{h}\mathbf{P}_{\gtrsim\lambda}F\|_{N}\lesssim\sum_{h=1}^{h_{0}}\sum_{N_{1}+N_{2}\leq N}l^{1-h-N_{1}}\|v\|_{1}\|F\|_{N_{2}+1}\lambda^{-h-1}$
@@
where $\mathcal{R}_{h}g$ has Fourier coefficient at frequency $k\in\mathbb{Z}^{3}\setminus\{0\}$ given by $\left[\nabla^{h}\mathcal{F}(\mathcal{R})\right](k)\mathcal{F}(g)(k)$, and the second-to-last inequality is due to Eq.˜A.42.
At last,
(A.57) $\|J_{2}\|_{N}$ $\lesssim\sum_{2^{j}\gtrsim\lambda}\sum_{k,\eta}|k|^{N}|\eta||\mathcal{F}(v_{l})(k-\eta)||\mathcal{F}(F_{j})(\eta)||k-\eta|^{h_{0}+1}|\eta-\sigma(\eta-k)|^{-h_{0}-2}$
$\lesssim\sum_{2^{j}\gtrsim\lambda}\sum_{k,\eta}|\eta|^{N-h_{0}-1}|\mathcal{F}(F_{j})(\eta)||k-\eta|^{h_{0}}|\mathcal{F}(\nabla v_{l})(k-\eta)|$
$\lesssim\sum_{2^{j}\gtrsim\lambda}\frac{l^{-h_{0}}}{2^{j(h_{0}+1)}}\sum_{|k|\leq 2^{j}}\|\nabla^{N}F_{j}\|_{L^{2}}\|\nabla v\|_{L^{2}}\lesssim(l\lambda)^{-h_{0}}\lambda^{2}\|v\|_{1}s_{0}\,a_{F}\lambda^{N}.$
where the second inequality is due to $l^{-1}<\frac{\lambda}{100}$ and $|k-\eta|\leq 2l^{-1}$, which imply that $|\eta-\sigma(\eta-k)|\geq\frac{|\eta|}{2}$ and $|k|\lesssim|\eta|$. All in all, Eqs.˜A.49, A.50, A.54, A.56 and A.57 prove the claim. ∎
@@
(A.58) $\sum_{h=1}^{h_{0}}\|\nabla^{h}v_{l}\nabla\mathcal{R}_{h}\mathbf{P}_{\gtrsim\lambda}F\|_{N}\lesssim\sum_{h=1}^{h_{0}}\sum_{N_{1}+N_{2}\leq N}l^{1-h}\|\nabla v_{l}\|_{N_{1}}\lambda^{-h}\|F\|_{N_{2}}.$
This inequality and Eq.˜A.57 imply that
@@
Under the same assumption of Propositions˜A.3 and A.4, it holds that
(A.60) $\|D_{t,l}\sum_{m,k}\epsilon^{\lambda}_{n_{0}}(k\cdot\xi_{m},a_{m,k})e^{i\lambda k\cdot\xi_{m}}\|_{0}\lesssim a_{F}(c^{-1}\vee\|v_{l}\|_{1})s_{n_{0}+1}\lambda^{-n_{0}-1},$
@@
For fixed $(m,k)$, we want to estimate $\|D_{t,l}\epsilon_{n_{0}}^{\lambda}\|_{0}$, given by Eq.˜A.22. we have to consider $\epsilon_{n_{0},j}(k\cdot\xi_{m},a_{m.k})$, which is given by Eq.˜A.9. Based on Eq.˜A.30 and $Z_{0}(r):=r^{-1}Z(r)$, we have that
(A.61) $\|D_{t,l}Z^{(n)}_{0}\|_{0}$ $\lesssim\|\int_{0}^{1}(1-s)(-s)^{n}[v_{l}(x)-v_{l}(x-yrs)]\cdot\nabla_{x}[(y\cdot\nabla)^{n+2}(k\cdot\xi_{m})(x-rsy)]ds\|_{0}$
$+\|\int_{0}^{1}(1-s)(-s)^{n}[D_{t,l}(y\cdot\nabla)^{n+2}]k\cdot\xi_{m}(x-rsy)ds\|_{0}$
@@
(A.62) $\|D_{t,l}(y\cdot\nabla)^{n}g\|_{0}\lesssim|y|^{n}\left[\|D_{t,l}g\|_{n}+\sum_{\begin{subarray}{c}m_{1}+m_{2}=n\vskip 0.1pt\\ m_{1}\geq 1\end{subarray}}\|v_{l}\|_{m_{1}}\|g\|_{m_{2}+1}\right].$
Then, similarly to Eqs.˜A.31 and A.32, we have that
@@
(A.66) $\|D_{t,l}e^{iZ(k\cdot\xi_{m})}\|_{0}$ $\lesssim\lambda\|\int_{0}^{1}(1-s)(y\otimes y):[(D_{t,l}\nabla^{2}(k\cdot\xi_{m})(x-ry)$
(A.67) $\quad+\left((v_{l}(x)-v_{l}(x-ry))\cdot\nabla\right)\nabla^{2}(k\cdot\xi_{m})(x-ry)ds]e^{i\lambda Z(r)}\|_{0}$
@@
Instead, Assumption 2 and Eq.˜A.62 imply that
@@
By following the computations used to get Eq.˜A.37, Eqs.˜A.64, A.66, A.65 and A.69 imply that

### v3

@@ A.1. Estimates for transport equations
We report some well-known estimates regarding smooth solutions $f:\mathbb{R}\times\mathbb{R}^{3}\to\mathbb{R}$ of the transport equation
(A.1) $\begin{cases}&\partial_{t}f+v\cdot\nabla f=g,\\ &f(t_{0},\cdot)=f_{0},\end{cases}$
where $v:\mathbb{R}\times\mathbb{R}^{3}\to\mathbb{R}^{3}$ is a smooth vector field and $g:\mathbb{R}\times\mathbb{R}^{3}\to\mathbb{R}$ is a smooth scalar field. We adopt the notation $f_{t}:=f(t,\cdot)$, and similarly for $v,g$, et cetera. The following proposition is a slight modification of [BuDLSzVi19, Proposition B.1].
@@
(A.5) $\|\nabla\phi_{t}-\mathrm{Id}\|_{0}$ $\lesssim|t-t_{0}|\|v_{[t_{0},t]}\|_{1},$
(A.6) $[\phi_{t}]_{N}$ $\lesssim|t-t_{0}|[v_{[t_{0},t]}]_{N},\quad\forall N\geq 2.$
@@ A.2. Antidivergence estimates
(A.7) $\mathscr{F}[T_{\mathfrak{m}}(f)](k)=\mathscr{F}[f](k)\mathfrak{m}(k),\quad k\in\mathbb{Z}^{3}.$
@@
Let $v\in C^{\infty}(\mathbb{T}^{3},\mathbb{R}^{3})$ and $l,\epsilon,\lambda$ be positive constants, and $\overline{n},n_{0}\in\mathbb{N}^{*}$ s.t. $\overline{n}\geq n_{0}+1$. Let $v_{l}:=\mathbf{P}_{\leq l^{-1}}v$ and consider, for any $m\in\mathbb{Z}$, the solution $\xi_{m}$ to the vector-valued transport equation
@@
Then, for every $N\leq\overline{n}$ and every $t\in\mathbb{R}$ we have
@@
with $\epsilon_{n_{0},j}$ being obtained by Lemma A.1 applied to the Fourier multiplier operator $\mathbf{P}_{2^{j}}$. Notice that, by construction, $\xi_{m}$ is smooth and space-periodic, since $v_{l}$ is so.
In order to show (A.20), let us consider $(m,k)\in\mathbb{Z}\times(\mathbb{Z}^{3}\setminus\{0\})$ and $j_{k}\in\mathbb{N}$ such that $2^{j_{k}}\geq\frac{5}{4}\lambda|k|$. Then
(A.23) $\mathbf{P}_{\gtrsim 3\lambda/8}=\sum_{3\lambda/8\leq 2^{s}\leq 2^{j_{k}}}\mathbf{P}_{2^{s}}+\mathbf{P}_{>2^{j_{k}}}=:T_{\mathfrak{m}_{1}}+\sum_{j>j_{k}}T_{\mathfrak{m}_{j}},$
@@
(A.24) $\operatorname{Supp}\,\mathfrak{m}_{1}\subset B(0,2^{j_{k}+1}),\,\quad\mathfrak{m}_{1}\equiv 1\,\text{ on }\,\overline{B(0,2^{j_{k}})}\setminus B\left(0,\frac{3}{4}\lambda|k|\right),\quad\mathfrak{m}_{j}\equiv 0\text{ on }\overline{B}\left(0,\frac{5}{4}\lambda|k|\right).$
Moreover for every $t\in\operatorname{Supp}_{t}a_{m,k}$, Assumption 1 of this proposition and (A.5) imply that
@@
Therefore, Lemma A.1 applied to $T_{m_{1}}$, and noticing that $\mathfrak{m}_{1}(\lambda\nabla(k\cdot\xi_{m}))\equiv 1$, $\nabla^{h}\mathfrak{m}_{1}(\lambda\nabla(k\cdot\xi_{m}))\equiv 0$ for every $h\geq 1$, giving:
(A.26) $T_{\mathfrak{m}_{1}}(a_{m,k}e^{i\lambda k\cdot\xi_{m}})$ $-\sum_{3\lambda/8\leq 2^{s}\leq 2^{j_{k}}}\mathcal{\epsilon}_{n_{0},s}(k\cdot\xi_{m},a_{m,k})e^{i\lambda k\cdot\xi_{m}}$
@@
At last, (A.20) follows from applying Lemma A.1 to $T_{\mathfrak{m}_{j}}$ for all $j>j_{k}$. Indeed,
@@
Now, for fixed $(m,k)$, we want to estimate $\|\epsilon_{n_{0}}^{\lambda}(k\cdot\xi_{m},a_{m,k})\|_{0}$. Based on (A.22), we have to consider $\epsilon_{n_{0},j}(k\cdot\xi_{m},a_{m.k})$, which is given by (A.9).
Let $Z$ be defined by (A.12) and $Z_{0}(r):=r^{-1}Z(r)$, then
@@
Therefore, Assumption 1 and Proposition A.1 imply that
@@
Then, by means of (A.13), (A.14), we have that
@@
Then, (A.32), (A.34) imply that
@@
So, (A.20) allows to conclude as follows:
@@
(A.41) $\lesssim a_{F}[s_{0}\,\lambda^{N-1}+s_{n_{0}+1}\lambda^{-n_{0}-1}\lambda^{N}],$
@@
Under the same assumptions of Proposition A.2, let $l>\frac{100}{\lambda}$ and $w\in C^{\infty}(\mathbb{T}^{3},\mathbb{R}^{3})$ such that there exists a non-decreasing sequence $(b_{n})_{n\in\mathbb{N}}$ satisfying for every $n,m\in\mathbb{N}$
@@
where $D_{t,l}:=\partial_{t}+v_{l}\cdot\nabla$ and $j\leq{\overline{n}}$, and that for any $n\in\mathbb{N}$
@@
Then for every $N\leq{\overline{n}-2}$ and $h_{0}\in\mathbb{N}\setminus\{0,1\}$ it holds
(A.46) $\|D_{t}\mathcal{R}F(t)\|_{N}$ $\lesssim\Big[s_{0}(c\lambda)^{-1}+c^{-1}s_{n_{0}+1}\lambda^{-n_{0}-1}+s_{0}\lambda^{-1}\|v\|_{1}+\|v\|_{0}\lambda^{-n_{0}}s_{n_{0}+1}$
(A.47) $\quad+(1+\lambda^{-n_{0}}s_{n_{0}+1})(b_{0}+l\|v\|_{1})+s_{0}(l\lambda)^{-h_{0}}\lambda^{2}\|v\|_{1}\Big]\lambda^{N}a_{F},$
@@
Since $D_{t,l}F=\sum_{k,m}D_{t,l}a_{m,k}e^{i\lambda\xi_{m}\cdot k}$, then the previous Proposition A.2 implies that
@@
$\lesssim\lambda^{N+1}\|v\|_{0}\left\|\sum_{m,k}\epsilon^{\lambda}_{n_{0}}(k\cdot\xi_{m},a_{m,k})\right\|_{0}$
@@
where the last inequality is due to (A.38). To study $I_{1}$, we preliminarily observe that by [RT25, Proposition 4.12] the operator $\mathcal{R}$ can be seen as a Fourier multiplier with coefficients $\mathscr{R}(k)$, $k\in\mathbb{Z}^{3}\setminus\{0\}$, and since $I_{1}$ has zero space average we can rewrite
(A.55) $-I_{1}$ $=\sum_{\begin{subarray}{c}k,\eta\in\mathbb{Z}^{3}\setminus\{0\}\end{subarray}}i\eta e^{ik\cdot x}\mathscr{F}[v_{l}](k-\eta)\mathscr{F}[\mathbf{P}_{\gtrsim\lambda}F](\eta)\left(\mathscr{R}(k)-\mathscr{R}(\eta)\right).$
We can extend the function $\mathscr{R}$ to a smooth function defined on $\mathbb{R}^{3}$, still denoted by $\mathscr{R}$ with a slight abuse of notation, that coincides with the expression given in [RT25, Equation 4.17] for every $k\in\mathbb{R}^{3}\setminus B(0,1/2)$. Since $l^{-1}<\frac{\lambda}{100}$ by assumption and by construction $\mathscr{F}[\mathbf{P}_{\gtrsim\lambda}F](\eta)=0$ for $\eta\lesssim\lambda$, $\mathscr{F}[v_{l}](k-\eta)=0$ for $|k-\eta|>2l^{-1}$, for every $k,\eta\in\mathbb{Z}^{3}\setminus\{0\}$ contributing to the sum above it holds $|\eta-\sigma(\eta-k)|\geq\frac{|\eta|}{2}\gtrsim\lambda$ for every $\sigma\in[0,1]$. As a consequence, we can apply integration by parts and the mean value theorem to obtain:
(A.56) $-I_{1}$ $=\sum_{\begin{subarray}{c}k,\eta\in\mathbb{Z}^{3}\setminus\{0\}\end{subarray}}i\eta e^{ik\cdot x}\mathscr{F}[v_{l}](k-\eta)\mathscr{F}[\mathbf{P}_{\gtrsim\lambda}F](\eta)\left(\mathscr{R}(k)-\mathscr{R}(\eta)\right)$
$=\sum_{\begin{subarray}{c}k,\eta\in\mathbb{Z}^{3}\setminus\{0\},\\ |\eta|\gtrsim\lambda\end{subarray}}i\eta e^{ik\cdot x}\mathscr{F}[v_{l}](k-\eta)\mathscr{F}[F](\eta)\Bigg(\sum_{h=1}^{h_{0}}\frac{1}{h!}[(k-\eta)\cdot\nabla]^{h}\mathscr{R}(\eta)$
$+\frac{1}{h_{0}!}\int_{0}^{1}[(k-\eta)\cdot\nabla]^{h_{0}+1}\mathscr{R}(\eta+\sigma(k-\eta))(1-\sigma)^{h_{0}}d\sigma\Bigg)=:J_{1}+J_{2}.$
Let us observe that by [RT25, Equation 4.17] we have $|\nabla^{h}\mathscr{R}(\eta)|\sim|\eta|^{-1-h}$. Therefore
(A.57) $\|J_{1}\|_{N}$ $\lesssim\sum_{h=1}^{h_{0}}\sum_{N_{1}+N_{2}\leq N}l^{1-h-N_{1}}\|v\|_{1}\|F\|_{N_{2}+1}\lambda^{-h-1}$
@@
where the second-to-last inequality is due to (A.42). At last,
(A.58) $\|J_{2}\|_{N}$ $\lesssim\sum_{\begin{subarray}{c}k,\eta\in\mathbb{Z}^{3},\\ |\eta|\gtrsim\lambda\end{subarray}}|k|^{N}|\eta||\mathscr{F}[v_{l}](k-\eta)||\mathscr{F}[F](\eta)||k-\eta|^{h_{0}+1}|\eta-\sigma(\eta-k)|^{-h_{0}-2}$
$\lesssim\sum_{\begin{subarray}{c}k,\eta\in\mathbb{Z}^{3},\\ |\eta|\gtrsim\lambda\end{subarray}}|\eta|^{N-h_{0}-1}|\mathscr{F}[F](\eta)||k-\eta|^{h_{0}}|\mathscr{F}[\nabla v_{l}](k-\eta)|$
$\lesssim\frac{l^{-h_{0}}}{\lambda^{h_{0}+1}}\|\nabla^{N}\mathbf{P}_{\gtrsim\lambda}F\|_{L^{2}}\|\nabla v\|_{L^{2}}\lesssim(l\lambda)^{-h_{0}}\lambda^{2}\|v\|_{1}s_{0}\,a_{F}\lambda^{N}.$
where the second inequality is due to $|\eta-\sigma(\eta-k)|\geq\frac{|\eta|}{2}$ and $|k|\lesssim|\eta|$. All in all, (A.49), (A.50), (A.54), (A.57), (A.58) prove the claim. ∎
@@
(A.59) $\sum_{h=1}^{h_{0}}\|\nabla^{h}v_{l}\nabla^{h}\mathcal{R}\mathbf{P}_{\gtrsim\lambda}F\|_{N}\lesssim\sum_{h=1}^{h_{0}}\sum_{N_{1}+N_{2}\leq N}l^{1-h}\|\nabla v_{l}\|_{N_{1}}\lambda^{-h}\|F\|_{N_{2}}.$
This inequality and (A.58) imply that
@@
Under the same assumption of Proposition A.3, it holds that
(A.61) $\left\|D_{t,l}\sum_{m,k}\epsilon^{\lambda}_{n_{0}}(k\cdot\xi_{m},a_{m,k})e^{i\lambda k\cdot\xi_{m}}\right\|_{0}\lesssim a_{F}(c^{-1}\vee\|v_{l}\|_{1})s_{n_{0}+1}\lambda^{-n_{0}-1},$
@@
For fixed $(m,k)$, we want to estimate $\|D_{t,l}\epsilon_{n_{0}}^{\lambda}\|_{0}$, given by (A.22). we have to consider $\epsilon_{n_{0},j}(k\cdot\xi_{m},a_{m.k})$, which is given by (A.9). Based on (A.30) and $Z_{0}(r):=r^{-1}Z(r)$, we have that
(A.62) $\|D_{t,l}Z^{(n)}_{0}\|_{0}$ $\lesssim\left\|\int_{0}^{1}(1-s)(-s)^{n}[v_{l}(x)-v_{l}(x-yrs)]\cdot\nabla_{x}[(y\cdot\nabla)^{n+2}(k\cdot\xi_{m})(x-rsy)]ds\right\|_{0}$
$+\left\|\int_{0}^{1}(1-s)(-s)^{n}[D_{t,l}(y\cdot\nabla)^{n+2}]k\cdot\xi_{m}(x-rsy)ds\right\|_{0}$
@@
(A.63) $\|D_{t,l}(y\cdot\nabla)^{n}g\|_{0}\lesssim|y|^{n}\left(\|D_{t,l}g\|_{n}+\sum_{\begin{subarray}{c}m_{1}+m_{2}=n,\\ m_{1}\geq 1\end{subarray}}\|v_{l}\|_{m_{1}}\|g\|_{m_{2}+1}\right).$
Then, similarly to (A.31), (A.32), we have that
@@
(A.67) $\|D_{t,l}e^{iZ(k\cdot\xi_{m})}\|_{0}$ $\lesssim\lambda\Bigg\|\int_{0}^{1}(1-s)(y\otimes y):[(D_{t,l}\nabla^{2}(k\cdot\xi_{m})(x-ry)$
(A.68) $\quad+\left((v_{l}(x)-v_{l}(x-ry))\cdot\nabla\right)\nabla^{2}(k\cdot\xi_{m})(x-ry)ds]e^{i\lambda Z(r)}\Bigg\|_{0}$
@@
Instead, Assumption 2 and (A.63) imply that
@@
By following the computations used to get (A.37), (A.65), (A.67), (A.66), (A.70) imply that